# QUEST#4X: an extension of QUEST#4 for benchmarking multireference wavefunction methods

### v1

@@ 4.1 Reference data
Ideally, well-resolved experimental data can be taken as the reference. However, for the 24 radicals considered here, only some experimental 0-0 transitions280, 281, 282, 283, 284, 285, 286, 287, 288, 289, 290, 291, after correcting the geometric and vibrational effects179, 202, 283, are available for comparison (see Table 4.1). Nevertheless, they are enough to verify the accuracy of iCIPT2. It can be seen from Table 4.1 that iCIPT2, in conjunction with the AVTZ basis sets273, 274, 275, is indeed very accurate, except for the state of CH3 and the state of NO. It turns out that there exist substantial basis set effects for the two states145. When the AV5Z basis sets273, 274, 275 are used, the errors for the two states are reduced from 0.21 and 0.25 eV to 0.04 and 0.06 eV, respectively. Not surprisingly, iCIPT2 is very close to CIPSI (configuration interaction using a perturbative selection made iteratively)292, since both are near-exact methods. The deviations of the methods (UCC3, UCCSDT, UCCSDTQ, and CIPSI) adopted in Ref. 145 from iCIPT2 are further plotted in Fig. 2 for the 51 doublets reported therein. All these pinpoint that the iCIPT2/AVTZ VEEs for the 110 doublets and 39 quartets of the 24 radicals can be taken as a solid reference for calibrating other methods with the same basis sets.
[] Experimental and theoretical vertical excitation energies (VEEs in eV) of 26 doublet states Radical State Expt. iCIPT2/AVTZa CCSDT/AVTZb CIPSI/AVTZb CIPSI/CBSc C3H5 3.41 d 3.42 3.43 - - BeF 6.16 e 6.27 6.23 6.21 - 4.14 e 4.15 4.15 4.14 4.13 BeH 2.48 e 2.49 2.49 2.49 2.48 6.32 e 6.46 6.45 6.46 6.46 CH 3.94 e 3.98 4.03 3.98 3.96 2.88 e 2.91 2.94 2.91 2.90 3.26 f 3.29 3.31 3.29 3.28 CH3 5.73 e 5.86 (5.88) 5.86 5.85 5.88 7.44 e 7.65 (7.48) 7.65 7.65 7.48 CN 3.22 e 3.22 3.25 3.22 3.21 1.32 e 1.34 1.38 1.34 1.33 CNO 1.55 g 1.62 1.71 1.61 1.61 CO+ 5.82 e 5.80 5.70 5.81 5.80 3.26 e 3.27 3.26 3.28 3.26 F2BO 2.77 h 2.77 2.78 - - 0.70 h 0.70 0.71 - - F2BS 2.91 i 2.94 2.93 - - 0.46 i 0.49 0.48 - - H2PS 2.77 j 2.73 2.75 2.72 2.71 NCO 2.80 k 2.87 2.87 2.83 2.89 NO 5.92 e 6.11 (6.05) 6.13 6.13 6.12 7.03 e 7.28 (7.09) 7.29 - [7.21] OH 4.09 e 4.11 4.12 4.10 4.09 PH2 2.74 l 2.76 2.77 2.77 2.76 C2H3 3.22 m 3.28 3.31 3.26 - MDn 0.05 0.06 0.05 0.04 MADn 0.06 0.06 0.06 0.06 SDn 0.09 0.10 0.08 0.08 MAXn 0.25 0.26 0.21 0.18
@@
This work with AVTZ basis sets. In parentheses are AV5Z results. bRef. 145. cRef. 145. In square brackets is the CCSDTQ/AVQZ result. dRefs. 179, 280. eRefs. 202, 281, 282. fRefs. 283, 284. gRefs. 179, 285. hRefs. 179, 286. iRefs. 179, 287. jRefs. 179, 288. kRefs. 179, 289. lRefs. 179, 290. mRefs. 179, 291. nMD: mean deviation; MAD: mean absolute deviation; SD: standard deviation; MAX: maximum deviation. Unavailable data were excluded from the analysis.
@@ 4.2 QUEST#4X
Some criterion should first be established for adding additional excited states to the QUEST#4 dataset145. Since the highest VEE reported therein is 8.02 eV ( of OH), only those singly excited doublet and quartet states with VEEs up to 8.0 eV are to considered. To identify such states, iCIPT2/AVTZ calculations were performed on the low-lying excited states of the 24 radicals for each irreducible representation (irrep). Eventually, 59 doublet states (including 17 Rydberg states) and 39 quartet states (including 6 Rydberg states) were added to QUEST#4 (which has 51 doublet states, including 11 Rydberg states), leading to QUEST#4X. Although the AVTZ basis sets may not be sufficient for some of the Rydberg states, the CIPT2/AVTZ results can certainly be taken as reference for calibrating other methods with the same basis sets.
@@ 4.3 Active space
Unlike iCIPT2, the NEVPT2, SDSPT2, SDSCI, and ic-MRCISD calculations all require a proper choice of active spaces. The occupation numbers (ON) of the natural orbitals (NO) resulting from the iCIPT2 calculations were taken here as a guidance. Those NOs with ONs significantly different from zero or two were taken as the active orbitals, the number of which was, however, not allowed to exceed 14, in order to make the MR calculations feasible. The resulting active spaces for the 24 radicals are documented in Table 4.3. The SA-CASSCF calculations can be initiated and supervised by the iCAS approach (imposed automatic selection of complete active spaces)293, where once the valence atomic orbitals (see Table S2 in the Supporting Information) are prepared, all remaining steps are automated. In particular, the subspace matching algorithm therein preserves the characters of the orbitals (doubly occupied, active or virtual) throughout the self-consistent iterations.
@@ 4.4 Vertical Excitation Energies
The SS-NEVPT2, MS-NEVPT2, SDSPT2, SDSCI, and ic-MRCISD calculations on the VEEs of the 110 doublet and 39 quartet states of the 24 radicals in QUEST#4145 were performed with the above setup of active spaces. As a matter of fact, the SDSCI calculations yield SDSPT2 and SS/MS-NEVPT2 results for free. SDSCI also serves as the initial step of ic-MRCISD. The so-calculated VEEs are documented in Tables LABEL:table-ene. Apart from the overall error analysis shown in Table 4.4, the internal consistency of the methods is also analyzed in Table 4.4. Scatter plots for the iCIPT2 and X-TDA/SS-NEVPT2/MS-NEVPT2/SDSPT2/SDSCI/ic-MRCISD VEEs are further depicted in Fig. 3.
It can be seen from Fig. 3 that the three variants of MRPT2 (i.e., SS-NEVPT2, MS-NEVPT2, and SDSPT2) have a strong linear correlation with iCIPT2, with R2 larger than 0.995, which is companied by a mean absolute deviation (MAD) of ca. 0.1 eV. As can be seen from Table 4.4, the VEEs calculated by the three MRPT2 are closely aligned for more than 80% of the excited states, with the differences in between being less than 0.05 eV. However, SS-NEVPT2 fails occasionally. For instance, it fails to reproduce the ordering of the (6.83 eV) and (6.71 eV) states of H2PO. Despite some overestimates, both MS-NEVPT2 and SDSPT2 reproduce the correct ordering for the two states (MS-NEVPT2: 6.61 and 7.04 eV; SDSPT2: 6.64 and 7.05 eV), as compared to the iCIPT2 results (6.20 and 6.81 eV). A similar situation occurs also to ( and under ) and ( and under ) states of CH3. These represent examples where dynamic correlation revises significantly the coefficients of the reference functions, as shown in Table 4.4. Here, it deserves to be emphasized that, although MS-NEVPT2 and SDSPT2 are hardly distinguishable for the systems considered here (see Table 4.4), SDSPT2 does outperform MS-NEVPT2 for situations with multiple nearly degenerate states260.
[] Wavefunctions of excited states of H2PO and CH3 via CASSCF, MS-NEVPT2, and SDSPT2 H2PO Ground state: (10a’)1(3a”)2 (3a”)2 (11a’)0 (10a’)1 (4a”)1 72% 15% 18% 62% 3% 75% 87% 3% 7% 73% 84% 5% CH3 Ground state: (3a1)2(1b1)2(1b2)1 (1b2)1 to (5a1)0 (3a1)2 (1b2)1 96% 0% 0% 95% 20% 76% 75% 20% 42% 54% 54% 41% (1b2)1 to (2b1)0 (1b1)2 (1b2)1 96% 0% 0% 95% 20% 76% 75% 20% 48% 48% 48% 47%
Both SDSCI and ic-MRCISD have a perfect linear correlation with iCIPT2, with R2 being ca. 0.999, which is companied by a MAD of ca. 0.05 eV. SDSCI agrees with ic-MRCISD within 0.15 eV for all the excited states, given that the computational cost of SDSCI is merely that of one iteration of ic-MRCISD. More specifically, SDSCI agrees with ic-MRCISD within 0.05, 0.10, and 0.15 eV for 75%, 95%, and 100% of the excited states, respectively (cf. Table 4.4). In particular, this is not fortuitous, but holds also for other systems147. However, it deserves to be pointed out that both SDSCI and ic-MRCISD perform worse than the three MRPT2 for the state of nitromethyl (CH2NO2), see Table LABEL:table-ene.

### v2

@@ 4.1 Reference data
Ideally, well-resolved experimental data can be taken as the reference. However, for the 24 radicals considered here, only some experimental 0-0 transitions280, 281, 282, 283, 284, 285, 286, 287, 288, 289, 290, 291, after correcting the geometric and vibrational effects179, 202, 283, are available for comparison (see Table 4.1). Nevertheless, they are enough to verify the accuracy of iCIPT2. It can be seen from Table 4.1 that iCIPT2, in conjunction with the AVTZ basis sets273, 274, 275, is indeed very accurate, except for the state of CH3 and the state of NO. It turns out that there exist substantial basis set effects for the two states145. When the AV5Z basis set273, 274, 275 is used, the errors for the two states are reduced from 0.21 and 0.25 eV to 0.04 and 0.06 eV, respectively. Not surprisingly, the iCIPT2 results are very close to those by selected CI (sCI)145 since both methods are nearly exact. The deviations of the methods (UCC3, UCCSDT, UCCSDTQ, and sCI) adopted in Ref. 145 from iCIPT2 are further plotted in Fig. 2 for the 51 doublets reported therein. All these pinpoint that the iCIPT2/AVTZ VEEs for the 110 doublets and 39 quartets of the 24 radicals can be taken as a solid reference for calibrating other methods with the same basis sets.
[] Experimental and theoretical vertical excitation energies (VEEs in eV) of 26 doublet states Radical State Expt. iCIPT2/AVTZa 95% intervalb CCSDT/AVTZc sCI/AVTZa,d TBE/CBSe C3H5 3.41 f 3.40 0.02 0.03 3.43 - - BeF 6.16 g 6.27 0.00h 0.01 6.23 6.21 0.02 - 4.14 g 4.15 0.00 0.00 4.15 4.14 0.01 4.13 BeH 2.48 g 2.49 0.00 0.00 2.49 2.49 0.00 2.48 6.32 g 6.46 0.00 0.00 6.45 6.46 0.00 6.46 CH 3.94 g 3.98 0.00 0.00 4.03 3.98 0.00 3.96 2.88 g 2.91 0.00 0.00 2.94 2.91 0.00 2.90 3.26 i 3.29 0.00 0.00 3.31 3.29 0.00 3.28 CH3 5.73 g 5.86 0.00 (5.88 0.00) 0.00 ( 0.01) 5.86 5.85 0.01 5.88 7.44 g 7.65 0.00 (7.48 0.00) 0.00 ( 0.01) 7.65 7.65 0.01 7.48 CN 3.22 g 3.22 0.00 0.00 3.25 3.22 0.00 3.21 1.32 g 1.34 0.00 0.00 1.38 1.34 0.01 1.33 CNO 1.55 j 1.62 0.01 0.00 1.71 1.61 0.01 1.61 CO+ 5.82 g 5.80 0.00 0.00 5.70 5.81 0.00 5.80 3.26 g 3.27 0.00 0.00 3.26 3.28 0.00 3.26 F2BO 2.77 k 2.77 0.01 0.04 2.78 - - 0.70 k 0.70 0.04 0.03 0.71 - - F2BS 2.91 l 2.95 0.02 0.05 2.93 - - 0.46 l 0.50 0.00 0.04 0.48 - - H2PS 2.77 m 2.73 0.01 0.01 2.75 2.72 0.02 2.71 NCO 2.80 n 2.87 0.01 0.03 2.87 2.83 0.05 2.89 NO 5.92 g 6.11 0.01 (6.05 0.01) 0.01 ( 0.04) 6.13 6.13 0.02 6.12 7.03 g 7.28 0.01 (7.09 0.01) 0.00 ( 0.04) 7.29 - [7.21] OH 4.09 g 4.11 0.00 0.00 4.12 4.10 0.01 4.09 PH2 2.74 o 2.76 0.00 0.00 2.77 2.77 0.00 2.76 C2H3 3.22 p 3.28 0.00 0.01 3.31 3.26 0.02 - MDq 0.05 0.06 0.05 0.04 MADq 0.06 0.06 0.06 0.06 SDq 0.09 0.10 0.08 0.08 MAXq 0.25 0.26 0.21 0.18
@@
means extrapolation distance . In parentheses are the AV5Z results. bSum of the 95% confidence intervals of the ground and excited states. In parentheses are the AV5Z results. cRef. 145. dRef. 145. eRef. 145. sCI results calculated or corrected to at least AVQZ level. In brackets is the CCSDTQ/AVQZ result. fRefs. 179, 280. gRefs. 202, 281, 282. h6.22 0.00 eV with the AVTZ from Basis Set Exchange292. iRefs. 283, 284. jRefs. 179, 285. kRefs. 179, 286. lRefs. 179, 287. mRefs. 179, 288. nRefs. 179, 289. oRefs. 179, 290. pRefs. 179, 291. qMD: mean deviation; MAD: mean absolute deviation; SD: standard deviation; MAX: maximum deviation. Unavailable data were excluded from the analysis.
@@ 4.2 QUEST#4X
Some criteria should first be established for adding additional excited states to the QUEST#4 dataset145. Since the highest VEE reported therein is 8.02 eV ( of OH), only those singly excited doublet and quartet states with VEEs up to 8.0 eV are to be considered here. To identify such states, iCIPT2/AVTZ calculations were performed on the low-lying excited states of the 24 radicals for each irreducible representation (irrep). Eventually, 59 doublet states (including 17 Rydberg states) and 39 quartet states (including 6 Rydberg states) were added to QUEST#4 (which has 51 doublet states, including 11 Rydberg states), leading to QUEST#4X. Although the AVTZ basis sets may not be sufficient for some of the Rydberg states, the iCIPT2/AVTZ results can certainly be taken as the reference for calibrating other methods with the same basis sets.
@@ 4.3 Active space
Unlike iCIPT2, the NEVPT2, SDSPT2, SDSCI, and ic-MRCISD calculations all require a proper choice of active spaces. The occupation numbers (ON) of the natural orbitals (NO) resulting from the iCIPT2 calculations were taken here as guidance. Those NOs with ONs ranging from 1.98 to 0.02 were carefully analyzed as candidates for active orbitals. However, to make the MR calculations feasible, the number of active orbitals was kept below 15 (by deleting those virtual orbitals of the smallest ONs). The resulting active spaces for the 24 radicals are documented in Table 4.3. The SA-CASSCF calculations can be initiated and supervised by the iCAS approach (imposed automatic selection of complete active spaces)293, where once the valence atomic orbitals (see Table S2 in the Supporting Information) are prepared, all remaining steps are automated. In particular, the subspace matching algorithm therein preserves the characters of the orbitals (doubly occupied, active, or virtual) throughout the self-consistent iterations. The CASSCF energies are documented in Table S3 in the Supporting Information.
@@ 4.4 Vertical Excitation Energies
The SS-NEVPT2, MS-NEVPT2, SDSPT2, SDSCI, and ic-MRCISD calculations on the VEEs of the 110 doublet and 39 quartet states of the 24 radicals in QUEST#4145 were performed with the above setup of active spaces. The so-calculated VEEs are documented in Table LABEL:table-ene. Apart from the overall error analysis shown in Table 4.4, the internal consistency of the methods is also analyzed in Table 4.4. Scatter plots for the iCIPT2 and X-TDA/SS-NEVPT2/MS-NEVPT2/SDSPT2/SDSCI/ic-MRCISD VEEs are further depicted in Fig. 3.
It can be seen from Fig. 3 that the three variants of MRPT2 (i.e., SS-NEVPT2, MS-NEVPT2, and SDSPT2) have a strong linear correlation with iCIPT2, with R2 larger than 0.995, which is accompanied by a mean absolute deviation (MAD) of ca. 0.1 eV. As can be seen from Table 4.4, the VEEs calculated by the three MRPT2 are closely aligned for more than 80% of the excited states, with the differences in between being less than 0.05 eV. However, SS-NEVPT2 fails occasionally. For instance, it fails to reproduce the ordering of the (6.83 eV) and (6.71 eV) states of H2PO. Despite some overestimates, both MS-NEVPT2 and SDSPT2 reproduce the correct ordering for the two states (MS-NEVPT2: 6.61 and 7.04 eV; SDSPT2: 6.64 and 7.05 eV), as compared to the iCIPT2 results (6.20 and 6.81 eV). A similar situation occurs also to ( and under ) and ( and under ) states of CH3. These represent examples where dynamic correlation revises significantly the coefficients of the reference functions, as shown in Table 4.4. Here, it deserves to be emphasized that, although MS-NEVPT2 and SDSPT2 are hardly distinguishable for the systems considered here (see Table 4.4), SDSPT2 does outperform MS-NEVPT2 for situations with multiple near-degenerate states260.
Both SDSCI and ic-MRCISD have a perfect linear correlation with iCIPT2, with R2 being ca. 0.999, which is accompanied by a MAD of ca. 0.05 eV. SDSCI agrees with ic-MRCISD within 0.15 eV for all the excited states, given that the computational cost of SDSCI is merely that of one iteration of ic-MRCISD. More specifically, SDSCI agrees with ic-MRCISD within 0.05, 0.10, and 0.15 eV for 75%, 95%, and 100% of the excited states, respectively (cf. Table 4.4). In particular, this is not fortuitous but holds also for other systems147. However, it deserves to be pointed out that both SDSCI and ic-MRCISD perform worse than the three MRPT2 for the state of nitromethyl (CH2NO2), see Table LABEL:table-ene.
@@
[…]
As for the MRPT2 calculations, it is first noted that, although SS-NEVPT2 predicts correctly the relative ordering of the four states, it overestimates the excitation energies of the and states by 0.41 eV and 0.47 eV, respectively, but underestimates that of by 0.64 eV. To understand such failures of standard MRPT2, we analyzed the wavefunctions in detail. As can be seen from Table 4.4, the four states stem primarily from four transitions, two valence and two Rydberg transitions. For the state, which is predominantly a mixture of the two valence transitions, all MR methods yield similar results, with deviations from iCIPT2 by no more than 0.15 eV. In contrast, the wavefunctions of the other three states are more complex. At the CASSCF level, both and are dominated by Rydberg characters (>70%), while is more akin to a valence state yet mixed with 32% Rydberg character. The situation changes dramatically at the SDSCI and ic-MRCISD levels, where these states exhibit a strong mixing of all the four components, resulting in comparable Rydberg characters exceeding 50% for each state. Such revisions of the reference states cannot be captured by standard MRPT2 like SS-NEVPT2, but can be captured to some extents by their multi-state/quasi-degenerate variants. For instance, SDSPT2 reproduces essentially all the features of SDSCI/ic-MRCISD, whereas MS-NEVPT2 is much less successful, in the sense that its state contains only 25% Rydberg component, less than half of that of ic-MRCISD. Nevertheless, there is no significant difference in the overall accuracy of MS-NEVPT2 and SDSPT2: compared with iCIPT2, MS-NEVPT2 overestimates the excitation energies of the to states by 0.07, 0.12, and 0.07 eV, respectively, while SDSPT2 exhibits deviations of 0.22, 0.15, and -0.07 eV, respectively. It appears that MS-CASPT2 also performs well for these states298.